OpenStudy (anonymous):

Which equation is equivalent to 3x – 4y = 8?

OpenStudy (anonymous):

where are the options??

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

Y=3/4x+2

OpenStudy (anonymous):

that's A.

OpenStudy (anonymous):

B. Y=-3/4x+2

OpenStudy (anonymous):

C. Y=-3/4x-2

OpenStudy (anonymous):

D. Y=3/2-2

OpenStudy (yttrium):

It is C.

OpenStudy (anonymous):

thanks

OpenStudy (phi):

if you start with 3x – 4y = 8 add +4y to both sides: 3x -4y + 4y = 8+4y simplify -4y +4y to 0 to get 3x = 8 + 4y add -8 to both sides 3x - 8 = 8+ -8 + 4y simplify 8 + -8 to 0 to get 3x -8 = 4y or 4y= 3x-8 divide both sides and *all* terms by 4 4y/4 = 3/4 x - 8/4 simplify y = (3/4) x -2